OpenStudy (anonymous):

Identify the curve for the parametric equation of x=-2+cos(t) and y=sin(t)

OpenStudy (anonymous):

The curve is a circle with center in (-2,0) and radius = 1

OpenStudy (anonymous):

Thank you !!! how do you do it ?

OpenStudy (anonymous):

The general parametric equations of a circle is: x=a+r*cos(t), y=b+r*sin(t), where (a,b) is the center of circle, and radius = r